Solve. –3(10x – 3) + 10(3x – 3) = x – 7
Answer by fractalier(6550)

From
–3(10x – 3) + 10(3x – 3) = x – 7
let us expand, collect like terms and solve...we get
-30x + 9 + 30x - 30 = x - 7
-21 = x - 7
-14 = x
x = -14